OpenStudy (anonymous):

3x + 3y - 2z = 13 6x + 2y - 5z = 13 7x + 5y - 3z = 26 Solve for x,y,and z. Show your work. @Mathematics

jhonyy9 (jhonyy9):

multiply fisrt by -2 and assume with second -6x-6y+4z=-26 6x+2y-5z=13 ---------------- 0 -4y-z=-13 multiply second by -7and 3rd by 6 -42x-14y+35z=-91 42x+30y-18z=156 -------------------- 0 16y+17z= 65 so than we have -4y-z=-13 divide both sides by -1 and will get 4y+z=13 so than z=13-4y 16y+17z=65 - if z=13-4y than 16y+17(13-4y)=65 16y+221-68y=65 -52y=-156 divide by -1 52y=156 y=156/52 =78/26 =39/13 = 3 so y=3 than -4(3)-z=-13 so -12-z=-13 so -z= -1 so z=1 than we get from 3x+3y-2z=13 so 3x+9-2=13 so 3x=6 so x=2 than x=2 y=3 z=1

OpenStudy (anonymous):

Thanks can u help me with 3 more?